OpenStudy (anonymous):

find the domain for each function: a.f(x)=5/x+2 b.f(x)=√3-x

OpenStudy (anonymous):

do you know what domain is?

OpenStudy (freckles):

That first one is a fraction. You just need to think what x can be and that is the domain. What x cannot be is not apart of the domain. Fractions don't exist when the bottom is 0 <--so start here to find the domain of the first fraction

OpenStudy (freckles):

For example 5x/(x+3) does not exist at x=-3 because that would make the (x+3) , 0. So the domain of 5x/(x+3) is all real numbers except x=-3 To say that an interval notation you say \[(- \infty , -3) \cup (-3,infty)\]